Graph the following function:

y= ex / x2

Ensure that you include the following properties:

  • Asymptote, x, y – intercepts, Domain and Range
  • First derivative, Maximum and Minimum
  • Second derivative, inflection points
  • Regions or increasing and decreasing
  • Regions or concave up and concave down
  • Sketch of the graph
1
Expert's answer
2022-04-05T14:14:29-0400

Assimptotes x=0, y=0

It is no x,y- interceptes.

"Domain: \n(\n\u2212\n\u221e\n,\n0\n)\n\u222a\n(\n0\n,\n\u221e\n)\n,\n{\nx\n|\nx\n\u2260\n0\n}"

"Range: \n(\n0\n,\n\u221e\n)\n,\n{\ny\n|\ny\n>\n0\n}"

"y'=\frac{e^x(x-2)}{x^3}"

"\frac{e^x(x-2)}{x^3}=0" , x=2, "y=e^2\/4"


"y"=-\frac{4xe^x-x^2e^x-6e^6}{x^4}"

At x=2 y">0, so

(2,e2/4) -local minimum

No inflection points, because "y" \not=0"

"Increasing\ on: 
(
−
∞
,
0
)
,
(
2
,
∞
)"

"Decreasing\ on: 
(
0
,
2
)"

"Concave\ up\ on 
(
−
∞
,
0
)
 since 
f
''
(
x
)
 is\ positive"

"Concave\ up\ on 
(
0
,
∞
)
 since 
f
''
(
x
)
 is\ positive"